OpenStudy (anonymous):

Integrate: x^2 sinx^3

OpenStudy (anonymous):

hi lana

OpenStudy (anonymous):

the answer at the back is -cosx^3/3

OpenStudy (lalaly):

i forgot the - and sorry du= 3x^2

OpenStudy (lalaly):

u=x^3 du=3x^2 dx now substitute integration sinu/2 du = 1/3 cosu +c substitute x back =- 1/3 cos x^3 +C

OpenStudy (anonymous):

thanks guys:)

OpenStudy (lalaly):

hey adrian:D